OpenStudy (anonymous):

Solve for x a) 8^5 =32^3x

OpenStudy (solomonzelman):

3x is the exponent, right ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 8^5 =32^{3x} }\) \(\LARGE\color{blue}{ \bf (2^{3})^5 =32^{3x} }\) \(\LARGE\color{blue}{ \bf (2^5)^{3} =32^{3x} }\) \(\LARGE\color{blue}{ \bf 32^{3} =32^{3x} }\) can you do it now ?

OpenStudy (solomonzelman):

I really though the logarithms were necessary, but oh well....

OpenStudy (anonymous):

how would i solve for x

OpenStudy (solomonzelman):

If you do not know how to solve for x after I (not even "pretty much, but JUST) just gave you the answer, then no matter how I am sorry to say this, but I really can't help you -:(

OpenStudy (solomonzelman):

I am giving you a chance. \(\LARGE\color{blue}{ \bf 2^{3}=2^{b} }\) THEREFORE \(\LARGE\color{blue}{ \bf 3=b }\)

OpenStudy (anonymous):

so i add the 3's that would make x = 6

OpenStudy (solomonzelman):

No, lost case, I am out from here !

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I would teach you, but you are probably not willing.

OpenStudy (anonymous):

i am

OpenStudy (solomonzelman):

OK

OpenStudy (solomonzelman):

You will need to be engaging a lot, to show the willingness. \(\LARGE\color{blue}{ \bf 9=9 }\) right? got it? Now, \(\LARGE\color{blue}{ \bf 9=3^{2} }\) correct ?

OpenStudy (anonymous):

yes i understand that

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 3^{2}=3^{2} }\) correct?

OpenStudy (anonymous):

ya they are the same

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 3^{b}=3^{2} }\) these THREEs have to have equal exponenets to be equal to each other, so therefore we can conclude that \(\LARGE\color{blue}{ \bf b=2 }\) RIGHT ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 32^{3}=32^{3x}~~~. }\) What can you say about this ?

OpenStudy (anonymous):

something has to equal to 32^3

OpenStudy (anonymous):

we sub in a number for x

OpenStudy (solomonzelman):

the THIRTY-TWOs have to be equal to each other, and so do their exponents.

OpenStudy (anonymous):

ok 32^3(1)

OpenStudy (anonymous):

the number wont change because the 1

OpenStudy (solomonzelman):

yes, so x= ?

OpenStudy (anonymous):

x= 32768

OpenStudy (anonymous):

am i right?

OpenStudy (solomonzelman):

OMG, I am stopping this question, even though I am not a moderator !!! \(\normalsize\color{black}{ \bf ▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄▄ }\)

OpenStudy (anonymous):

why?

OpenStudy (solomonzelman):

Because you need to talk to your teacher, and get help from a teacher or a special tutor. And I am not joking, this is very very serious !!

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Yes, please do... for your own sake.

OpenStudy (anonymous):

ok